OpenStudy (anonymous):

I3x+14/xI=4 find the solution set.

OpenStudy (amoodarya):

the eq. has no real root but 1st step solve 3x+14/x=4 multiply by x 3x^2+14=4x has no root 2st step solve 4x+14/x=-4 like that has no root